# TwinMarket: A Scalable Behavioral and Social Simulation for Financial Markets

### v2

@@ D.3 Belief
The belief structure is modeled along five key dimensions: the users’ view on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment. These dimensions capture a comprehensive range of factors that influence investor beliefs, forming the basis for their desires and ultimately their investment intentions. These align with the key components of the Belief, Desire, and Intention (BDI) framework, which posits that behavior is a result of these three interconnected factors.
@@
D.4 Rumor
@@
Figure 16 and Figure 17 exemplify the news inputs used in our study. Specifically, Figure 16 illustrates factual reports from reputable news sources, representing a typical information environment an investor might encounter. In contrast, Figure 17 showcases fabricated or exaggerated negative news snippets designed to simulate the spread of misinformation, which we refer to as rumors.
The deliberate introduction of these fabricated rumors is not merely an academic exercise; it’s a critical step that distinguishes our research. We are not content with modeling investor behavior under ideal conditions of accurate information. Instead, we are confronting the very real, often chaotic, information landscape in which investors actually operate. By injecting these carefully crafted rumors, we are directly examining the vulnerability of retail investors to the insidious effects of misinformation, unveiling the subtle mechanisms by which market panic can be manufactured and disseminated. This innovative approach allows us to not only observe the direct impact of negative rumors on trading behavior but also to model how these waves of manipulated information cascade through investment communities, potentially leading to systemic instability.
@@
D.5 Trading System
We have designed an order matching trading system based on the A-share market’s call auction rules. The system follows the principles of “price priority, time priority" and implements price limit restrictions (±10%). The system matches orders based on the quoted prices and order volumes of both buyers and sellers, using the maximum transaction volume principle. It also tracks large orders and capital flows, while updating users’ positions. This system provides TwinMarket with a trading environment that closely resembles real market conditions.
Appendix E More Validation
E.1 Mirco-level Validation
Figure [18, 19, 20, 21, 22, 23, 24] show that the users we have constructed, guided by prompts, autonomously exhibit both rational investment strategies and irrational behavioral biases during decision-making, mirroring real-world investors.
E.2 Scaling up Validation
This section presents the scale-up validation results as part of the appendix.
Table 8: Platform Statistics for Scaled-Up Validation
Metric Value
RMSE 0.02
MAE 0.02
Correlation 0.77
Estimated Lag 0 time units
Max Cross-correlation 0.9998
In order to evaluate the robustness and generalizability of our scale-up approach, we performed a series of evaluation methods. The detailed evaluation metrics for these experiments are summarized in Table 8.
The Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) both measure the magnitude of the errors in our model’s predictions; in our case both the RMSE and MAE are 0.02, indicating a low average prediction error, which means that our model predictions are closely aligned with actual market movements. The correlation with the real index is measured as the Pearson Correlation Coefficient, which is 0.77 in our experiments, indicating a strong positive correlation between our model’s predictions and the real market index. This high correlation demonstrates that our model is capable of capturing the main trends of real market behavior, and shows good potential for real-world applications. Further, the Estimated Lag between our predictions and the actual market movements is 0 time units, meaning that our model is able to react to market movements in a timely manner. Finally, the Max Cross-correlation of 0.9998 means that the predictions of our model aligns very closely with the direction of the actual price movements, with almost a 1:1 correspondence between when the model expects a jump and when the jump actually happens.
These results demonstrate that our model performs well at larger scales, and can reliably simulate investor behavior, while accurately mirroring general market trends. The combination of low prediction errors, strong positive correlation with real indexes, zero lag, and high cross correlation highlights the validity of our approach, and the ability of the model to provide an accurate market simulation. These results further provide a solid foundation for subsequent research.
Figure 18: Example of Fundamental Analysis.
Figure 19: Example of Technical Analysis.
Figure 20: Examples of Overconfidencce Bias.
Figure 21: Examples of Loss Aversion and Disposition Effect.
Figure 22: Examples of Lottery Preference.
Figure 23: Examples of Insufficient Diversification.
Figure 24: Examples of Herd Behavior.
Appendix F Prompt
Figures 25 to 36 show the overall process of our TwinMarket.

### v3

@@ G.1 Agent Belief Structure and Representation
The belief score ranges from 1 to 5 and is computed as the average of five key dimensions: the users’ view on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment at each update step. These dimensions capture a comprehensive range of factors that influence investor beliefs, forming the basis for their desires and ultimately their investment intentions. These align with the key components of the Belief, Desire, and Intention (BDI) framework, which posits that behavior is a result of these three interconnected factors.
@@
G.2 Social Interaction Dynamics
G.2.1 Rationale for Social Network Edge Formulation
As detailed in Section 3.3, the formulation of edges within TwinMarket’s simulated social network models users exhibiting greater similarity in trading behaviors with stronger interpersonal connections. This represents a principled design choice, adopted to facilitate the study of emergent social phenomena while acknowledging the inherent challenge of precisely replicating the proprietary and often opaque recommendation algorithms employed by real-world social media platforms.
The primary assumption underpinning this edge construction—that correlated trading patterns often signify underlying shared information channels or mutual influence—is supported by established findings [85], which demonstrated that instances of correlated trading among investors frequently arise from common information linkages. This empirical evidence provides a robust justification for utilizing trading behavior similarity as a proxy for the potential strength of social connections pertinent to information dissemination and interpersonal influence within a financial market context.
Furthermore, it is imperative to note that the principal research objective of TwinMarket is the investigation of emergent collective behaviors, rather than the exact algorithmic replication of any specific existing social platform. To this end, the adoption of well-established network modeling principles, notably the concept of homophily (i.e., the propensity of individuals to associate and bond with similar others) based on observable characteristics like trading patterns, offers a tractable and theoretically sound foundation for our simulation.
G.2.2 Rationale for Hot Score Mechanism Design
Accurately modeling the dynamic influence of information propagation and content relevance in financial social environments is a notoriously challenging task. Commercial social media platforms treat their specific content ranking and recommendation algorithms as closely guarded proprietary information, making direct replication infeasible. To address this within TwinMarket, we designed a “hot score” mechanism (with its formula detailed in Section 3.3) for dynamically ranking posts. This mechanism is inspired by established principles from the wide use of time decay factor inrecommendation systems [86, 87], and is tailored to reflect the unique characteristics and temporal dynamics of information flow on social media platforms within a financial context [88].
By integrating the principled time decay with user-driven feedback, the hot score mechanism provides a transparent and interpretable dynamic ranking system. It prioritizes content that is timely, deemed relevant by the community, and actively engaging, thereby shaping the information landscape to which agents are exposed in a realistic manner.
G.2.3 Post-propagation Chain Dynamics
The visibility and influence of posts within TwinMarket’s social environment are shaped by their propagation dynamics. When updating a post’s score, both the breadth (number of direct and indirect shares/interactions) and depth (length of the sharing chain) of its propagation are considered. Specifically, if a user interacts with an original (root) post by reposting or liking any of its subsequently propagated versions, this engagement contributes to the cumulative score of the initial root post. This mechanism ensures that content generating significant and sustained engagement continues to gain prominence and visibility as it diffuses across the network.
Figure 29 provides a visual example of such post propagation, illustrating how a single initial post can give rise to multiple distinct propagation chains. Each branch in these chains represents a unique path through which the information, or the post itself, spreads from user to user. This branching structure reinforces the overall reach and potential influence of widely shared content.
Figure 29: An illustrative example of a post-propagation chain, where an initial post (root) branches out through user interactions (e.g., reposts), demonstrating the spread of information.
G.3 Simulation of Misinformation: Rumor Injection
@@
The news inputs utilized in our study are exemplified in Figure 30 and Figure 31. Figure 30 presents examples of factual reports from reputable news sources, simulating the information environment typically encountered by investors. Conversely, Figure 31 displays examples of fabricated or exaggerated negative news snippets, hereafter referred to as rumors, which are designed to simulate the propagation of misinformation.
The deliberate introduction of these fabricated rumors constitutes a critical methodological step that distinguishes our research. Our approach moves beyond modeling investor behavior solely under idealized conditions of veridical information. Instead, it directly confronts the complex and often chaotic information landscape characteristic of real-world financial markets. Through the injection of these carefully crafted rumors, our study directly examines the vulnerability of investors to the deleterious effects of misinformation. This allows for an investigation into the subtle mechanisms through which market panic can be instigated and propagated.
@@
G.4 Order-Driven Trading System Design
To ensure a realistic and dynamic market environment within TwinMarket, we have implemented a order-driven trading system based on the A-share market’s call auction rules.
Our system operates based on the fundamental principles of “price priority, time priority”, which means that buy orders with higher prices and sell orders with lower prices take precedence. Among orders at the same price level, those submitted earlier are prioritized. Furthermore, to maintain market stability and realism, the system incorporates price limit restrictions ( ±$\pm 10\%$ from the previous closing price), preventing extreme, unrealistic price swings within a single trading session.
The core matching logic adheres to the maximum transaction volume principle. During a matching cycle, the system identifies a price at which the largest possible volume of shares can be traded by matching compatible buy and sell orders. It also meticulously tracks significant orders (“large orders”) and aggregate capital flows, providing insights into potential market-moving activities. Upon successful matching, trades are executed, and the system instantaneously updates the positions and cash balances of the participating agents.
By implementing this detailed order-driven trading system, TwinMarket creates a transactional environment that closely mirrors the operational intricacies of actual financial markets. This not only enhances the realism of the simulated trading experience for the agents but also ensures that emergent market phenomena, such as price volatility and liquidity fluctuations, are grounded in plausible micro-level interactions.
G.5 Management of Potential Data Leaks
A critical consideration when using Large Language Models (LLMs) for simulating historical or ongoing scenarios is the potential for information leakage from the LLM’s vast training data, which might include knowledge of real-world events or stock behaviors pertinent to the simulation period. Similarly, inherent biases within the LLM could unintentionally influence agent behavior. TwinMarket incorporates several safeguards to mitigate these risks and ensure that agent decisions are driven by the simulated environment and their defined personas, rather than external, pre-existing knowledge.
Our primary strategies to prevent data leakage and minimize the impact of LLM biases include:
  1. 1.
    Temporal Abstraction and Relative Time: To prevent the LLM from anchoring the simulation to specific real-world events contemporary to our chosen simulation timeframe, we abstract absolute time references. Within the simulation, agents are not exposed to absolute year markers like “2023”. Instead, they perceive time relatively, such as “Day N of the simulation”, “June 15th”, or through the sequence of events and information presented. This makes it significantly harder for the LLM to leverage its training data about specific occurrences during the actual 2023 period that might coincide with the simulation’s internal timeline. All historical data (news, announcements, stock prices) provided during initialization or as inputs are presented within this relative or abstracted time context.
  2. 2.
    Entity Anonymization with Neutral Identifiers: Specific names of real-world stocks, indices, and potentially prominent companies or individuals mentioned in news/announcements are systematically replaced with neutral, anonymized identifiers (e.g., Index TLEI, Stock CSP as shown in Table 10). This de-identification process severs direct links between the simulated entities and their real-world counterparts that the LLM might have knowledge about. By operating with these generic labels, agents’ decisions are based on the characteristics and information flow associated with these anonymized entities within the simulation, rather than any pre-conceived notions or historical performance data the LLM might associate with the actual underlying real-world entities.
  3. 3.
    Validation through Ablation Studies: To assess the impact of potential information leakage, we conducted ablation experiments, as detailed in Section 6 and Appendix D.2. Initial observations suggest that agents’ decision-making is more strongly influenced by the immediate interactions and emergent dynamics within the simulation (e.g., social posts from other agents, recent simulated price movements) rather than an over-reliance on the historical news or announcements provided at the start or during the simulation when making trading decisions. This indicates that the structured information environment and agent interaction mechanisms may effectively direct the LLM’s focus inward, reducing the influence of potentially leaked knowledge.
By implementing these measures, we aim to create a controlled environment where emergent behaviors are a product of the simulation’s design and agent interactions, rather than an artifact of the LLM’s existing knowledge base.
Appendix H Prompt
Figures 32 to 43 show the overall process of our TwinMarket.